OpenStudy (anonymous):

how would I factor 2b^4-32?

OpenStudy (anonymous):

2(b4-16) =2(b^4 - 2^4) =2(b^2-2^2)(b^2+2^2) =2(b- sqrt2)(b+ sqrt2)(b^2+2^2)

OpenStudy (anonymous):

First factor out the 2: \[2(b^4-16)\] Then, note that it is in the form of \[a^2-b^2=(a-b)(a+b)\] \[2(b^4-16)=2[ (b^2)^2-4^2 ]\]\[=2(b^2+4)(b^2-4)\]

OpenStudy (anonymous):

thanks